OpenStudy (anonymous):

Pls help solve the equation (w+3)^-1/5=1/4

OpenStudy (anonymous):

oops got it w=1021

OpenStudy (anonymous):

(-1/5)log(w+3) = log(1/4) log(w+3)/5 = log(4) w+3=(10 OR e)^(5log(4))